Minden Company introduced a new product last year for which it is trying to find an optimal seling price. Marketing studies suggest that the company can increase sales by 5,000 units for each $2 reduction in the selling price. The companys present selling price is $99 per unit, and variable expenses are $89 per unit Fixed expenses are $838,200 per year, The present annual sales volume (at the $99 selling price) is 25,100 units Required: 1. What is the present yearly operating income or loss? S163,400 2. What is the present break-even point in unit sales and in dollar sales? Braak-evan point in units 27,933 Break-even point in dolar sales 3. Assuming that the marketing studies are correct, what is the maximum annual profit that the company can earn? At how many units and at what selling price per unit would the company generate this proft? Maximum profit1,002,000 Number of units Selling price 50.100 85 4-a. What would be the break-even point in unit sales and in dollar sales using the selling price you determined in (3) above (e.g., the selling price at the level of maximum profits)? Break-even point in units 41,900

Minden Company introduced a new product last year for which it is trying to find an optimal selling price. Marketing studies suggest that the company can increase sales by 5,000 units for each $2 reduction in the selling price. The company’s present selling price is $99 per unit, and variable expenses are $69 per unit. Fixed expenses are $838,200 per year. The present annual sales volume (at the $99 selling price) is 25,100 units.

Required:

1. What is the present yearly operating income or loss?

2. What is the present break-even point in unit sales and in dollar sales?

3. Assuming that the marketing studies are correct, what is the maximum annual profit that the company can earn? At how many units and at what selling price per unit would the company generate this profit?

4-a. What would be the break-even point in unit sales and in dollar sales using the selling price you determined in (3) above (e.g., the selling price at the level of maximum profits)?
Answer #1

1. present yearly operating income/loss:

sales...................................................$2484900

Less: variable cost...............................1731900

contribution margin................753000

Less: fixed cost.....................................838200

net loss...................................(85200)

2.break even point:

p/v ratio=contribution margin/sales

=753000/2484900

=30.30%

In break even point, contribution margin=fixed cost

then,

sales=838200/30.30%

=$2766337   

sales units=2766337/99

=27943 units

3.maximum profit earned , if company opt the suggestion

contribution margin per unit sales unit contribution margin

    20 40100 802000

22 45100 992200

20 50100 1002000

18    55100 991800

the company can earned maximum profit =contribution margin -fixed cost

=1002200-838200

=174000

selling price=$20 per unit

sales unit =50100 units

4. break even point os sales if opt suggestion:

Break-even sales unit=fixed cost/ contribution margin per unit

=838200/20

=41910 units

Break-even sales=41910*89

=$3729990